OpenStudy (anonymous):

If sinӨ = -1/4 and Ө terminates in the third quadrant, find the exact value of sin2Ө

jimthompson5910 (jim_thompson5910):

You need to use the formula sin(2x) = 2*sin(x)*cos(x) so you'll need to find the value of cos(theta) first

OpenStudy (anonymous):

cos(theta) = 1

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

Put it in the calc O_O

jimthompson5910 (jim_thompson5910):

put what in the calc? lol

jimthompson5910 (jim_thompson5910):

cos(1/4) ?

OpenStudy (anonymous):

Oh no and it's -1/4 not 1/4 right?

OpenStudy (anonymous):

and that is .999990480

jimthompson5910 (jim_thompson5910):

oh yeah, but you don't do cos(-1/4) either

OpenStudy (anonymous):

It comes out to the same thing

jimthompson5910 (jim_thompson5910):

you need to solve the following for x sin^2 + cos^2 = 1 (-1/4)^2 + x^2 = 1 1/16 + x^2 = 1 .... ... x = ??

jimthompson5910 (jim_thompson5910):

once you get that, you'll know what cos(theta) is

OpenStudy (anonymous):

oh my bad >_<

jimthompson5910 (jim_thompson5910):

that's fine

OpenStudy (anonymous):

x=sqrt of 15/16 = .9682458366 Did I do that right?

jimthompson5910 (jim_thompson5910):

better, I would keep the exact form

jimthompson5910 (jim_thompson5910):

since we're in Q3, cosine is negative, so cos(theta) = -sqrt(15/16) sin(2theta) = 2*sin(theta)*cos(theta) sin(2theta) = 2*(-sqrt(15/16))*(-1/4) sin(2theta) = 2*(-sqrt(15)/sqrt(16))*(-1/4) sin(2theta) = 2*(-sqrt(15)/4)*(-1/4) sin(2theta) = sqrt(15)/8

OpenStudy (anonymous):

Thank you c:

jimthompson5910 (jim_thompson5910):

np